Media services charges $30.00 for a phone and $20.00/month for its economy plan. The equation c=20t+30 describes the total cost, c, of operating a Media services phone for t months. The total cost for 8 months of service is how much? If a customer has only $80.00 available, how many months of service can she receive? Please help. Thank you.
Answer by stanbon(75887)

Media services charges $30.00 for a phone and $20.00/month for its economy plan.
The equation c=20t+30 describes the total cost, c, of operating a Media services phone for t months.
-------------------------------
The total cost for 8 months of service is how much?
C(8) = 20*8+30 = 190
-----------------------------------------
If a customer has only $80.00 available, how many months of service can she receive?
Solve 20t + 30 = 80
20t = 50
t = 2.5 months.
======================
Cheers,
Stan H.
========================